# Stochastic Optimal Control of an Industrial Power-to-Heat System with High-Temperature Heat Pump and Thermal Energy Storage

### v2

@@ 3.2 State Constraints
Due to the configuration of the P2H system, the TES temperature is bounded, i.e. $R_{n}\in[r_{\text{min}},r_{\text{max}}]$ for all $n$. More precisely, ensuring constant SG inlet and outlet temperatures by fluid bypass regulation implies that the storage temperature $R_{n}$ cannot be greater than $r_{\text{max}}=T^{\text{SG,in}}$ and cannot fall below $r_{\text{min}}=T^{\text{SG,out}}$.
In contrast, the exogenous state variables are generally based on our modeling approach in (16). Wind speeds are by nature non-negative and potentially unbounded, implying that $W_{n}\in(0,\infty)$. Unlike wind speeds, the electricity prices are also allowed to become negative and we have $S_{n}\in(-\infty,\infty)$. A negative price may occur in the case of overproduction of electricity and at the same time a lower demand in the grid. In addition, in this case, producers are penalized for feeding in additional power, while consumers are rewarded for using electrical energy from the grid.
@@ 3.3 Control Constraints
The various state and operational constraints mentioned in the previous subsections imply constraints on the control and lead to state-dependent sets of feasible controls from which the controller can select the actions. In particular, the control $A_{n}$ for the period $[t_{n},t_{n+1})$ can only be selected such that technical upper limits for the HTHX outlet temperature $T^{\text{HT,out}}_{n}=\tau^{\text{HT,out}}_{\text{max}}$ and the HTHX inlet temperature $T^{\text{HT,in}}_{n}=\tau^{\text{HT,in}}_{\text{max}}$ are not exceeded. While the maximal HTHX outlet temperature $\tau^{\text{HT,out}}_{\text{max}}$ is directly related to the maximal compressor shaft speed, the maximal temperature $\tau^{\text{HT,in}}_{\text{max}}$ is set by system constraints of the HTHP. Furthermore, the controller needs to consider the time-varying charging and discharging factors $l^{\text{C}},l^{\text{D}}$, which determine the bypasses, can only take values in $[0,1]$ and must ensure that the TES temperature $R$, which is also time-varying, does not leave the range $[r_{\text{min}},r_{\text{max}}]=[T^{\text{SG,out}},T^{\text{SG,in}}]$. Given that the state at the beginning of the period $[t_{n},t_{n+1})$ is $X_{n}=x=(r,w,s)$, the control $A_{n}$ can be selected from a set of feasible controls of the form
@@
where $\underline{a}(r),\overline{a}(r)$ are piecewise linear functions of the TES temperature, which we derive in (196) and (199) in Appendix G. Figure 3 illustrates the derived set of feasible controls and their dependence on $R$ for the system parameters listed in Appendix C.
Figure 3: Visualization of the control constraints (28) as a function of the TES temperature. Upper bound $\overline{a}(r)$ (blue), lower bound $\underline{a}(r)$ (green), and the sets of feasible controls for the control $A_{n}(x)$ (red). If the TES is almost full, the upper bound $\overline{a}(r)$ is decreasing and approaches zero to prevent overheating during charging. If the TES is not sufficiently full, the decreasing lower limit $\underline{a}(r)$ prevents undercooling during discharging. In both cases, the heat flow is throttled accordingly. The maximum of the positive upper bound $\overline{a}(r)$ and the minimum of the negative lower bound $\underline{a}(r)$ result from the maximal inlet and outlet temperatures of the HTHX, respectively.
@@ 3.4 Operational Costs
In our model, we consider the running operational costs $C_{n}:\pazocal{X}\times\pazocal{A}\to\mathbb{R}$ in each of the periods $n=0,\ldots,N-1$. These are defined as the expected cumulated costs in the period $[t_{n},t_{n+1})$, given that at time $t_{n}$ the state $X_{n}=x=(r,w,s)^{\top}$ and the control $A_{n}=a$ is chosen, and read as
@@
$\Psi(t,W(t),S(t),a)=S(t)(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{+}-\delta S_{sell}(t)(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{-},$ (30)
with $\pi^{\text{H}}$ introduced in (14). The conditional expectation $\mathbb{E}_{n,x,a}(\cdot)=\mathbb{E}(\cdot|X_{n}=x,~{}A_{n}=a)$ emphasizes the dependence on the current time grid point $t_{n}$ and the current state $X_{n}=x$ as well as the action $A_{n}=a$ selected at this state. Further, we denote by $z^{+}=\max(z,0)$ and $z^{-}=\max(-z,0)$ the positive and negative part of $z\in\mathbb{R}$, respectively. The functional $\Psi$ in (30) is divided into two parts: (i) the costs for buying electricity from the grid at price $S(t)$ and (ii) the revenue for selling excess energy at a lower price $S_{sell}(t)$. Here, $\delta\in\{0,1\}$ indicates whether selling is allowed or not. If it is not allowed to sell excess energy to generate revenue, we set $\delta=0$. In this case, the surplus or overproduction of energy is discarded. For $\delta=1$ energy is fed into the grid for reduced market price $S_{sell}$ given by
$S_{sell}(t)=S(t)-\eta(t),$ (31)
where $\eta:[0,t_{\text{E}}]\to\mathbb{R}^{+}$ is called spread. This spread reflects transaction fees, taxes or the willingness of the grid operator to buy energy only at a certain discount on the market price $S(t)$. A special situation occurs in times of negative market prices $S$, which are often caused by energy overproduction. In this case, buying from the grid leads to a reward, i.e. one gets paid for purchasing energy. Selling, on the other hand, causes additional cost to keep the grid stable due to the abundance of energy. Here, a spread $\eta>0$ leads to a further reduction of the selling price, which results in higher costs for feeding energy into the grid and therefore makes selling less attractive. For more details on the computation of the running costs $C_{n}(x,a)$ and in particular the conditional expectation in (29), see Appendix H.1.
@@
where the functions $g_{\text{Pen}}:\pazocal{X}\to\mathbb{R}^{+}$ and $g_{\text{Liq}}:\pazocal{X}\to\mathbb{R}^{-}$ describe the amount of thermal energy required to adjust the TES temperature from $r$ to $r_{\text{crit}}$. In the case of penalization, $g_{\text{Pen}}$ units of thermal energy must be fed in, while for liquidation, $g_{\text{Liq}}$ units are withdrawn. If the TES is not sufficiently filled, i.e. $r<r_{\text{crit}}$, a penalty applies for energy consumption at a fixed price $s_{\text{Pen}}\geq 0$, depending on the respective temperature difference. In the opposite case, excess energy in the TES is liquidated, which means that energy is sold at the fixed price $s_{\text{Liq}}\geq 0$, which generates a revenue that appears as a negative terminal cost $G_{N}$. It should be noted that this definition of the terminal cost includes a worthless expiration of the TES by setting $s_{\text{Pen}}=s_{\text{Liq}}=0$.
@@ 3.6 Transition Operator
The transition from one state to another within the feasible set $\pazocal{X}$ is mathematically described by the transition operator. For state $X_{n}$, action $A_{n}=u_{n}(X_{n})$ at time point $t_{n}$ and a random disturbance $Z_{n+1}=(Z^{W}_{n+1},Z^{S}_{n+1})\sim\pazocal{N}(0_{2},\textrm{I}_{2})$, the state dynamics of the system is defined by the transition operator as
@@
$g^{R}(x,a)=r+\frac{1}{m_{\text{S}}c_{\text{S}}}a\Delta t,\quad\text{for }x=(r,w,s).$ (37)
@@
For the discrete-time system we will consider a filtered probability space with the filtration $\mathbb{F}=(\pazocal{F}_{n})_{n=0,...,N}$ where the $\sigma$-algebras $\pazocal{F}_{n}=\sigma(Z_{1},...,Z_{n})$ are generated by the independent random variables $Z_{1},\dots,Z_{n}$, and $\pazocal{F}_{0}=\{\emptyset,\Omega\}$ is the trivial $\sigma$-algebra.
@@ 3.7 Performance Criterion and Optimization Problem
with the running and terminal cost defined in Subsection 3.4. We denote $\pazocal{U}$ as the set of all admissible policies $u$ such that the objective function (41) is well-defined and $A_{n}=u_{n}(X_{n})$ satisfies the control constraints for all $n=0,\ldots,N-1$. An admissible policy $u\in\pazocal{U}$ is called optimal if
@@
In practice, it is not tractable to minimize over the space of all admissible policies $\pazocal{U}$. We will see that the performance criterion (41) satisfies a recursion property from which we are able to deduce an alternative way for obtaining the value function. The following theorems can be found in Bäuerle and Rieder (Baeuerle2011, , p. 21-23). Theorem 3.4 states that the objective function (41) satisfies a recursion property.
@@
Using Theorem 3.4, we obtain that the value function $V_{n}(x)$ is the solution of the Bellman equation.
@@
The Bellman equation reduces the problem of finding an optimal policy $u^{*}\in\pazocal{U}$ to a recursion in which only the optimal actions for each time point $n$ have to be found.
@@
Many applications and algorithms use an alternative performance criterion for $J^{u}_{n}$, which is often called the state action function and denoted by $Q_{n}^{u}$ Bertsekas1996 . It is defined for all $(x,a)\in\pazocal{X}\times\pazocal{A}_{n}(x)$ as
@@
Note that the expectation is conditional to $X_{n}=x,u_{n}(x)=a$, and given a policy $u$ with deterministic decision rules $u_{n}$ it holds that
$J_{n}^{u}(x)=Q^{u}_{n}(x,u_{n}(x)),$
for all $n=0,\ldots,N$. As the name suggests, the function $Q^{u}_{n}$ assigns a value to each feasible state action pair $(x,a)$ instead of a single value for each state $x$ as is the case with the performance criterion $J^{u}_{n}$. By Theorem 3.4 $Q_{n}^{u}$, satisfies the recursion
$\begin{split}Q_{n}^{u}(x,a)&=C_{n}(x,a)+\mathbb{E}_{n,x,a}[J_{n+1}^{u}(X_{n+1})]\\ &=C_{n}(x,a)+\mathbb{E}_{n,x,a}[Q_{n+1}^{u}(X_{n+1},u_{n+1}(X_{n+1}))].\end{split}$ (46)
@@
with optimal policy $u^{*}=(u_{n}^{*})_{n=0,\ldots,N-1}$ and solves the following Q-version of the Bellman equation annurev-statistics-031219-041220
@@ 4 Backward Dynamic Programming
In this section, we introduce an algorithm that solves the stochastic optimal control problem. This method is inspired by the Bellman equation (44) and is known as backward dynamic programming (BDP). We will also discuss issues that naturally arise when using BDP and how to address them.
@@ 4.1 Backward Recursion Algorithm
Note that given a state $X_{n}=x$ and action $A_{n}=a$, the next state $X_{n+1}$ is obtained by the transition operator (36), which depends on the random disturbance $Z_{n+1}$. Furthermore, the randomness in the subsequent state is only induced by the disturbance allowing to take the expectation with respect to the random variable $Z_{n+1}$, which is independent of $X_{n}$ and $A_{n}$, instead of $X_{n+1}$ giving
@@
Algorithm 1 summarizes the backward recursion procedure of BDP and is used to solve the MDP derived in Section 3.
@@
Step 1: Initialize $V_{N}(x)=G_{N}(x)$ for all $x\in\pazocal{X}$ and set $n=N-1$
Step 2: Compute for all $x\in\pazocal{X}$ the value function at time $n$
@@
$u_{n}^{*}(x)\in\underset{a\in\pazocal{A}_{n}(x)}{\operatorname*{arg\,min}}\bigg{\{}C_{n}(x,a)+\mathbb{E}[V_{n+1}(\pazocal{T}_{n}(x,a,Z_{n+1}))]\bigg{\}}$ (51)
Step 3: If $n>0$ set $n=n-1$ and go to step 1 else stop the algorithm.
@@ 4.2 Approximate Solution of the Bellman Equation
BDP may face some issues. For instance, if the state and action space are large, it may suffer from the curse of dimensionality, or in the case of continuous spaces, the optimization problem in Step 2 of Algorithm 1 may be difficult to solve. Another practical issue is the calculation of the expected value for all $x\in\pazocal{X}$, which may be computationally intractable due to the unavailability of closed-form expressions. Since it may be difficult to solve the Bellman equation exactly, we need to make certain simplifications to solve the issues described above.
@@
Firstly, the state space $\pazocal{X}\subset\mathbb{R}^{3}$ given in (33) is discretized into distinct grid points. The value function is then calculated and saved for the given reference grid points. For the discretization, the seasonalities $\mu_{W}$ and $\mu_{S}$ of the Ornstein-Uhlenbeck processes (16), are used to construct time-varying sets of grid points. The advantage of introducing this time dependence is that the value function is only calculated for regions of interests, i.e. subsets of $\pazocal{X}$ that are more likely to appear at certain times. This leads to the family of discretized state spaces
$\widetilde{\pazocal{X}}_{n}=\{r_{1},\ldots,r_{n_{\text{R}}}\}\times\{w_{1,n},\ldots,w_{n_{\text{W}},n}\}\times\{s_{1,n},\ldots,s_{n_{\text{S}},n}\}\qquad n=0,\ldots,N-1,$ (52)
with $n_{R},n_{W},n_{S}\in\mathbb{N}$. The specific choice of grid points used in our calculations can be found in Appendix I. Note that the discretization for the TES temperature is the same at every time point and only those of wind speed and electricity price change. The discrete structure of the state space $\widetilde{\pazocal{X}}_{n}$ allows to solve the Bellman equation for each $x\in\widetilde{\pazocal{X}}_{n}$ separately. To solve the problem of minimization over the action space $\pazocal{A}_{n}(x)\subset\mathbb{R}^{2}$ in BDP, the action space $\pazocal{A}_{n}(x)$ is also discretized into grid points $\widetilde{\pazocal{A}}_{n}(x)=\{a_{1},\ldots,a_{n_{\text{A}}}\}$ for given $n_{A}\in\mathbb{N}$. Therefore, the minimization consists of calculating a value for each action and picking the action with the smallest value.
@@
Given a state $X_{n}=x$ and action $A_{n}=a$, the conditional expected value in the Bellman equation with respect to the next state $X_{n+1}$ is given by unconditional expectation with respect to the random disturbance $Z_{n+1}$, see (50). Let $\pazocal{Z}=\{z_{1},\ldots,z_{L}\}$ be a set of values of $Z_{n+1}$ and denote $\widehat{Z}_{n+1}$ as the discrete random variable taking values in $\pazocal{Z}$. Further let $p_{l}=\mathbb{P}(\widehat{Z}_{n+1}=z_{l})$ be the corresponding probability that $\widehat{Z}_{n+1}$ takes value $z_{l}$, then the expected value can be approximated as weighted sum
@@
The set $\pazocal{Z}$ is called a quantizer of $Z_{n+1}$ and defines a partition on $\mathbb{R}^{2}$ into $L$ subsets $C(z_{l}),~{}l=1,\ldots,L$, where each point $z_{l}$ is uniquely assigned to a subset. As a consequence the probability $p_{l}$ corresponds to the probability that $Z_{n+1}$ takes values in $C(z_{l})$, i.e.
$p_{l}=\mathbb{P}(Z_{n+1}\in C(z_{l})).$ (54)
The calculation of these probabilities often requires solving high-dimensional integrals over the subsets $C(z_{l})$ with respect to the density of $Z_{n+1}$, which may become computationally infeasible as the dimension of the random disturbance grows.
@@
A practical problem is caused by the quantizer $\pazocal{Z}$ of $Z_{n+1}$. In general, the states $x_{l}=\pazocal{T}_{n}(x,a,z_{l})$ do not coincide with the grid point $\widetilde{\pazocal{X}}_{n+1}$ for which the value function $V_{n+1}$ is calculated and saved. If the points are allocated in between existing values, an interpolation can be used to determine $V_{n+1}(x_{l})$, otherwise this value must be calculated by extrapolation. In this paper, we use linear interpolation if $x_{l}=\pazocal{T}_{n}(x,a,z_{l})$ is in between existing point and extrapolate with the value of the nearest neighbor in the set $\widetilde{\pazocal{X}}_{n+1}$. The corresponding extrapolation errors are usually larger than those resulting from interpolation. In any case, the value $V_{n+1}(x_{l})$ is weighted by the probability $p_{l}$, and if this it is small, the corresponding error introduced becomes will be smaller. For this reason, the probabilities $p_{l}$ can be used to reduce and control the errors in the calculation of the value function. Let us also note that an appropriate choice of the discretized state space $\widetilde{\pazocal{X}}_{n+1}$ helps to mitigate extrapolation. Details on the construction of the discretization, which takes this fact into account, can be found in Appendix I.
@@
Due to the approximation of the expected value and action space discretization, the value function obtained is an approximation of the exact one, denoted by $\widetilde{V}$. Therefore, the control corresponding to $\widetilde{V}$ is an approximation of the optimal control. The calculation of the value function $\widetilde{V}_{n}$ for $x\in\widetilde{\pazocal{X}}_{n}$ in the BDP scheme 1 reduces to

### v3

@@ 3.2 State Constraints
Due to the configuration of the P2H system, the TES temperature is bounded, i.e., $R_{n}\in[r_{\text{min}},r_{\text{max}}]$ for all $n$. More precisely, ensuring constant SG inlet and outlet temperatures by fluid bypass regulation implies that the storage temperature $R_{n}$ cannot be greater than $r_{\text{max}}=T^{\text{SG,in}}$ and cannot fall below $r_{\text{min}}=T^{\text{SG,out}}$.
In contrast, the exogenous state variables are generally based on our modeling approach in (16). Wind speeds are by nature non-negative and potentially unbounded, implying that $W_{n}\in(0,\infty)$. Unlike wind speeds, the electricity prices are also allowed to become negative and we have $S_{n}\in(-\infty,\infty)$. A negative price may occur in the case of overproduction of electricity, while at the same time there is a lower demand in the grid. In addition, in this case, producers are penalized for feeding in additional power, while consumers are rewarded for using electrical energy from the grid.
Figure 3: Visualization of the control constraints (27) as a function of the TES temperature. Upper bound $\overline{a}(r)$ (blue), lower bound $\underline{a}(r)$ (green), and the sets of feasible controls for the control $A_{n}(x)$ (red). If the TES is almost full, the upper bound $\overline{a}(r)$ is decreasing and approaches zero to prevent overheating during charging. If the TES is not sufficiently full, the decreasing lower limit $\underline{a}(r)$ prevents undercooling during discharging. In both cases, the heat flow is throttled accordingly. The maximum of the positive upper bound $\overline{a}(r)$ and the minimum of the negative lower bound $\underline{a}(r)$ result from the maximal inlet and outlet temperatures of the HTHX.
@@ 3.3 Control Constraints
The various state and operational constraints mentioned in the previous subsections imply constraints on the control and lead to state-dependent sets of feasible controls from which the controller can select the actions. In particular, the control $A_{n}$ for the period $t_{n},t_{n+1})$ can only be selected such that technical upper limits for the HTHX outlet temperature $T^{\text{HT,out}}_{n}=\tau^{\text{HT,out}}_{\text{max}}$ and the HTHX inlet temperature $T^{\text{HT,in}}_{n}=\tau^{\text{HT,in}}_{\text{max}}$ are not exceeded. While the maximal HTHX outlet temperature $\tau^{\text{HT,out}}_{\text{max}}$ is directly related to the maximal compressor shaft speed, the maximal inlet temperature $\tau^{\text{HT,in}}_{\text{max}}$ is set by system constraints of the HTHP. Furthermore, the controller needs to consider that the time-varying charging and discharging factors $l^{\text{C}},l^{\text{D}}$, which determine the bypasses, can only take values in $[0,1]$ and must ensure that the TES temperature $R$, which is also time-varying, does not leave the range $r_{\text{min}},r_{\text{max}}]=[T^{\text{SG,out}},T^{\text{SG,in}}]$. Given that the state at the beginning of the period $t_{n},t_{n+1})$ is $X_{n}=x=(r,w,s)$, the control $A_{n}$ can be selected from a set of feasible controls
@@
where $\underline{a}(r),\overline{a}(r)$ are piecewise linear functions of the TES temperature, which we derive in (93) and (96). Figure 3 illustrates the derived set of feasible controls and their dependence on $R$ for the system parameters listed in the Supplementary Material B.
@@ 3.4 Operational Costs
In our model, we consider the running operational costs $C_{n}:\pazocal{X}\times\pazocal{A}\to\mathbb{R}$ in each of the periods $n=0,\ldots,N-1$. These are defined as the expected cumulated costs in the period $t_{n},t_{n+1})$, given that at time $t_{n}$ the state $X_{n}=x=(r,w,s)^{\top}$ and the control $A_{n}=a$ is chosen, and read as
@@
$\Psi(t,W(t),S(t),a)=S(t)(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{+}-\zeta S_{\text{sell}}(t)(\pi^{\text{H}}(a)-P^{\text{W}}(t))^{-},$ (29)
with $\pi^{\text{H}}$ introduced in (14). The conditional expectation $\mathbb{E}_{n,x,a}(\cdot)=\mathbb{E}(\cdot|X_{n}=x,~A_{n}=a)$ emphasizes the dependence on the current time grid point $t_{n}$ and the current state $X_{n}=x$ as well as the action $A_{n}=a$ selected at this state. Further, we denote by $z^{+}=\max(z,0)$ and $z^{-}=\max(-z,0)$ the positive and negative part of $z\in\mathbb{R}$, respectively. The functional $\Psi$ in (29) is divided into two parts: (i) the costs for buying electricity from the grid at price $S(t)$ and (ii) the revenue for selling excess energy at a lower price $S_{\text{sell}}(t)$. Here, $\zeta\in\{0,1\}$ is a user defined model parameter that indicates if selling is allowed or not. If it is not allowed to sell excess energy to generate revenue, we set $\zeta=0$. In this case, the surplus or overproduction of energy is discarded. For $\zeta=1$, energy is fed into the grid for a reduced market price $S_{\text{sell}}$ given by
$S_{\text{sell}}(t)=S(t)-\eta(t),$ (30)
where $\eta:[0,t_{\text{E}}]\to\mathbb{R}^{+}$ is called spread. This spread reflects transaction fees, taxes or the willingness of the grid operator to buy energy only at a certain discount on the market price $S(t)$. A special situation occurs in times of negative market prices $S$, which are often caused by energy overproduction. In this case, buying from the grid leads to a reward, i.e., one gets paid for purchasing energy. Selling, on the other hand, causes additional cost to keep the grid stable due to the abundance of energy. Here, a spread $\eta>0$ leads to a further reduction of the selling price, which results in higher costs for feeding energy into the grid and therefore makes selling less attractive. For more details on the computation of the running costs $C_{n}(x,a)$ and in particular the conditional expectation in (28), see Subsection Appendix H.1.
@@
where the functions $g_{\text{Pen}}:\pazocal{X}\to\mathbb{R}^{+}$ and $g_{\text{Liq}}:\pazocal{X}\to\mathbb{R}^{-}$ describe the amount of thermal energy required to adjust the TES temperature from $r$ to $r_{\text{crit}}$. In the case of penalization, $g_{\text{Pen}}$ units of thermal energy must be fed in, while for liquidation, $g_{\text{Liq}}$ units are withdrawn. If the TES is not sufficiently filled, i.e., $r<r_{\text{crit}}$, a penalty is applied for energy consumption at a fixed price $s_{\text{Pen}}\geq 0$, depending on the respective temperature difference. In the opposite case, excess energy in the TES is liquidated, which means that energy is sold at the fixed price $s_{\text{Liq}}\geq 0$, which generates a revenue that appears as a negative terminal cost $G_{N}$. It should be noted that this definition of the terminal cost includes a worthless expiration of the TES by setting $s_{\text{Pen}}=s_{\text{Liq}}=0$.
@@ 3.6 Transition Operator
The transition from one state to another, within the feasible set $\pazocal{X}$, is mathematically described by the transition operator. For state $X_{n}$, action $A_{n}=u_{n}(X_{n})$ at time point $t_{n}$ and a random disturbance $Z_{n+1}=(Z^{W}_{n+1},Z^{S}_{n+1})\sim\pazocal{N}(0_{2},\textrm{I}_{2})$, the state dynamics of the system is defined by the transition operator as
@@
$g^{R}(x,a)=r+\frac{1}{m_{\text{s}}c_{\text{S}}}a\Delta t,\quad\text{for}\quad x=(r,w,s).$ (37)
@@
For the discrete-time system we will consider a filtered probability space with the filtration $\mathbb{F}=(\pazocal{F}_{n})_{n=0,\dots,N}$ where the $\sigma$-algebras $\pazocal{F}_{n}=\sigma(Z_{1},\dots,Z_{n})$ are generated by the independent random variables $Z_{1},\dots,Z_{n}$, and $\pazocal{F}_{0}=\{\emptyset,\Omega\}$ is the trivial $\sigma$-algebra.
@@ 3.7 Performance Criterion and Optimization Problem
with the running and terminal cost defined in Subsection 3.4. We denote by $\pazocal{U}$ as the set of all admissible policies $u$ such that the objective function (41) is well-defined and $A_{n}=u_{n}(X_{n})$ satisfies the control constraints for all $n=0,\ldots,N-1$. An admissible policy $u\in\pazocal{U}$ is called optimal if
@@
In practice, it is not tractable to minimize over the space of all admissible policies $\pazocal{U}$. We will see that the performance criterion (41) satisfies a recursion property from which we are able to deduce an alternative way for obtaining the value function. The following theorems can be found in Bäuerle and Rieder (Baeuerle2011, , p. 21-23). Theorem 3.4 states that the objective function (41) fulfils a recursion property.
@@
Using Theorem 3.4, we obtain that the value function $V_{n}(x)$ is the solution of the well-known Bellman equation.
@@
The Bellman equation reduces the problem of finding an optimal policy $u^{*}\in\pazocal{U}$ to a recursion in which only the optimal actions for each time point $n$ must be found.
@@
Many applications and algorithms use an alternative performance criterion for $J^{u}_{n}$, which is often called the state action function and is denoted by $Q_{n}^{u}$ Bertsekas1996 . It is defined for all $(x,a)\in\pazocal{X}\times\pazocal{A}_{n}(x)$ as
@@
Note, that the expectation is conditional to $X_{n}=x,u_{n}(x)=a$, and given a policy $u$ with deterministic decision rules $u_{n}$ it holds that
$J_{n}^{u}(x)=Q^{u}_{n}(x,u_{n}(x)),$ (46)
for all $n=0,\ldots,N$. As the name suggests, the function $Q^{u}_{n}$ assigns a value to each feasible state action pair $(x,a)$, instead of a single value for each state $x$ as is the case with the performance criterion $J^{u}_{n}$. By Theorem 3.4, we get that $Q_{n}^{u}$, satisfies the recursion
$Q_{n}^{u}(x,a)$ $=C_{n}(x,a)+\mathbb{E}_{n,x,a}[J_{n+1}^{u}(X_{n+1})]$
$=C_{n}(x,a)+\mathbb{E}_{n,x,a}[Q_{n+1}^{u}(X_{n+1},u_{n+1}(X_{n+1}))].$ (47)
@@
with optimal policy $u^{*}=(u_{n}^{*})_{n=0,\ldots,N-1}$ and solves the Q-version of the Bellman equation annurev-statistics-031219-041220 below.
@@ 4 Backward Dynamic Programming
In this section, we introduce an algorithm that solves the stochastic optimal control problem. This method is inspired by the Bellman equation (44) and is known as backward dynamic programming (BDP). We will also discuss issues that naturally arise when using BDP and propose techniques on how to address them.
@@ 4.1 Backward Recursion Algorithm
Note that given a state $X_{n}=x$ and action $A_{n}=a$, the next state $X_{n+1}$ is obtained by the transition operator (36), which depends on the random disturbance $Z_{n+1}$. Furthermore, the randomness in the subsequent state is only induced by the disturbance allowing to take the expectation with respect to the random variable $Z_{n+1}$, instead of $X_{n+1}$ leading to
@@
Algorithm 1 summarizes the backward recursion procedure of BDP, which will be used to solve the MDP derived in Section 3.
@@
1: Initialize $V_{N}(x)=G_{N}(x)$ for all $x\in\pazocal{X}$ and set n=N1n=N-1
2: Compute for all $x\in\pazocal{X}$ the value function at time $n$
@@
$u_{n}^{*}(x)\in\underset{a\in\pazocal{A}_{n}(x)}{\operatorname*{arg\,min}}\bigg\{C_{n}(x,a)+\mathbb{E}[V_{n+1}(\pazocal{T}_{n}(x,a,Z_{n+1}))]\bigg\}$
3: If $n>0$ set n=n1n=n-1 and go to step 1 else stop the algorithm.
@@ 4.2 Approximate Solution of the Bellman Equation
BDP may face some issues. For instance, if the state and action space are large or high-dimensional, it may suffer from the curse of dimensionality, or in the case of continuous spaces, the optimization problem in step 2 of Algorithm 1 may be difficult to solve. Another practical issue is the calculation of the expected value for all $x\in\pazocal{X}$, which may be computationally intractable due to the unavailability of closed-form expressions. Since it may be difficult to solve the Bellman equation exactly, we need to make certain simplifications in order to solve the issues described above.
@@
Firstly, the state space $\pazocal{X}\subset\mathbb{R}^{3}$ given in (33) is discretized into distinct grid points. The value function is then calculated and saved for the given reference grid points. For the discretization, the seasonalities $\mu_{W}$ and $\mu_{S}$ of the Ornstein-Uhlenbeck processes (16), are used to construct time-varying sets of grid points. The advantage of introducing this time dependence is that the value function is only calculated for regions of interests, i.e., subsets of $\pazocal{X}$ that are more likely to appear at certain times. This leads to the family of discretized state spaces
$\widetilde{\pazocal{X}}_{n}=\{r_{1},\ldots,r_{n_{\text{R}}}\}\times\{w_{1,n},\ldots,w_{n_{\text{W}},n}\}\times\{s_{1,n},\ldots,s_{n_{\text{S}},n}\},\quad n=0,\ldots,N-1,$ (52)
with $n_{R},n_{W},n_{S}\in\mathbb{N}$. The specific choice of grid points used in our calculations can be found in Section Appendix I. Note, that the discretization points for the TES temperature is the same at every time point and only those of wind speed and electricity price change. The discrete structure of the state space $\widetilde{\pazocal{X}}_{n}$ allows us to solve the Bellman equation for each $x\in\widetilde{\pazocal{X}}_{n}$ separately. To solve the problem of minimization over the action space $\pazocal{A}_{n}(x)\subset\mathbb{R}^{2}$ in BDP, the action space $\pazocal{A}_{n}(x)$ is also discretized into grid points $\widetilde{\pazocal{A}}_{n}(x)=\{a_{1},\ldots,a_{n_{A}}\}$ for given $n_{A}\in\mathbb{N}$. Hence, the minimization consists of calculating a value for each action and picking the action with the smallest value.
@@
Given a state $X_{n}=x$ and action $A_{n}=a$, the conditional expected value in the Bellman equation with respect to the next state $X_{n+1}$ is given by an unconditional expectation with respect to the random disturbance $Z_{n+1}$, see (51). Let $\pazocal{Z}=\{z_{1},\ldots,z_{L}\}$ be a set of value that $Z_{n+1}$ can take and denote $\widehat{Z}_{n+1}$ as the discrete random variable taking values in $\pazocal{Z}$. Further let $p_{l}=\mathbb{P}(\widehat{Z}_{n+1}=z_{l})$ be the corresponding probability that $\widehat{Z}_{n+1}$ takes value $z_{l}$, then the expected value can be approximated as weighted sum
@@
The set $\pazocal{Z}$ is called a quantizer of $Z_{n+1}$ and defines a partition on $\mathbb{R}^{2}$ into $L$ subsets $\pazocal{C}(z_{l}),~l=1,\ldots,L$, where each point $z_{l}$ is uniquely assigned to a subset. As a consequence the probability $p_{l}$ corresponds to the probability that $Z_{n+1}$ takes values in $\pazocal{C}(z_{l})$, more precisely
$p_{l}=\mathbb{P}(Z_{n+1}\in\pazocal{C}(z_{l})).$ (54)
The calculation of these probabilities often requires solving high-dimensional integrals over the subsets $\pazocal{C}(z_{l})$ with respect to the density of $Z_{n+1}$. Below, we will explain how to numerically obtain these probabilities. But before we come to this, we want to point out a practical problem that is caused by the quantizer $\pazocal{Z}$.
@@
In general, the states $x_{l}=\pazocal{T}_{n}(x,a,z_{l})$ do not coincide with the grid points in $\widetilde{\pazocal{X}}_{n+1}$ for which the value function $V_{n+1}$ is calculated and saved. If the points are allocated in between existing values, an interpolation can be used to determine $V_{n+1}(x_{l})$, otherwise this value must be calculated by extrapolation. In this paper, we use linear interpolation if $x_{l}=\pazocal{T}_{n}(x,a,z_{l})$ is in between existing point and extrapolate with the value of the nearest neighbor in the set $\widetilde{\pazocal{X}}_{n+1}$. The corresponding extrapolation errors are usually larger than those resulting from interpolation. In any case, the value $V_{n+1}(x_{l})$ is weighted by the probability $p_{l}$, and if this is small, the corresponding error introduced will be scaled down by $p_{l}$. For this reason, the probabilities $p_{l}$ can be used to reduce and control the errors in the calculation of the value function. We also note that an appropriate choice of the discretized state space $\widetilde{\pazocal{X}}_{n+1}$ helps to mitigate extrapolation. Details on the construction of the discretization that takes this fact into account can be found in Section Appendix I.
@@
Due to the approximation of the expected value and action space discretization, the value function obtained is an approximation of $V$ and will be denoted by $\widetilde{V}$. Therefore, the control corresponding to $\widetilde{V}$ is an approximation of the optimal control. The calculation of the value function $\widetilde{V}_{n}$ for $x\in\widetilde{\pazocal{X}}_{n}$ in the BDP Algorithm 1 reduces to